OpenStudy (cucacula):

find the matrix A if : 2A - [1 0 -2]T = [2 0 [ 4 7 3 ] -3 4 0 8] anyone?

OpenStudy (anonymous):

First do the \(T\) operation.

OpenStudy (cucacula):

you mean transform the T?

OpenStudy (cucacula):

@wio

OpenStudy (anonymous):

It is transpose operation.

OpenStudy (anonymous):

do you know how to transpose a matrix?

OpenStudy (cucacula):

opss sorry typo . it would be like this? [ 4 1 7 0 3 -2 ] @wio

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

now add it to both sides.

OpenStudy (cucacula):

then?

OpenStudy (anonymous):

then multiply both sides by \(1/2\)

OpenStudy (cucacula):

can you give me example? @wio

OpenStudy (anonymous):

You don't know how to add matrices?

OpenStudy (anonymous):

This is no different than algebra. You're just using matrix operations instead.

OpenStudy (cucacula):

after i get like this [ 4 1 now i need to multiply to this matrix [2 0 7 0 -3 4 3 -2 ] 0 8]

OpenStudy (cucacula):

@wio

OpenStudy (anonymous):

No, you add it to both sides.

OpenStudy (cucacula):

okay so it would be like this right? [6 1 4 4 3 6] so the next step ? @wio

OpenStudy (anonymous):

Yes, so you have \[ 2A = \begin{bmatrix}6&1\\4&4\\3&6\end{bmatrix} \]

OpenStudy (anonymous):

Now you solve for \(A\) using scalar multiplication.

OpenStudy (cucacula):

okay,you mean like this : 2A = 2 . [ 6 1 [ 12 2 4 4 = 8 8 3 6 ] 6 12] like this ? @wio

OpenStudy (anonymous):

If you multiply both sides by \(2\), then \(2\times 2A = 4A\) and you still haven't isolated \(A\).

OpenStudy (cucacula):

so how to solve A using scalar multiplication? @wio

OpenStudy (anonymous):

Okay, we want \(c\) where \(c\times 2 = 1\). So what is \(c\)?

OpenStudy (cucacula):

1/2 @wio

OpenStudy (cucacula):

so now how to solve A? @wio

OpenStudy (anonymous):

Multiply both sides by \(1/2\).

OpenStudy (cucacula):

i get this = [ 3 1/2 2 2 3/2 3] is it true? @wio

OpenStudy (anonymous):

Yes

OpenStudy (cucacula):

then,next step? @wio

OpenStudy (anonymous):

You have found \(A\). What more do you want?

OpenStudy (cucacula):

so that's all? @wio

OpenStudy (cucacula):

thank you for helping @wio